OpenStudy (anonymous):

how to solve using the quadratic formula d^2-4d+1=0

OpenStudy (anonymous):

find two numbers whose sum = "-4" product = "+1"

OpenStudy (anonymous):

we cant.. so, we use the quadratic formula

OpenStudy (anonymous):

comparing with ax^2 +bx+c-0 sum of roots =-b/a =4 product of roots are c/a=1 hence roots are 1/4 and 4

OpenStudy (anonymous):

\[d=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

comparing with ax^2 +bx+c-0 sum of roots =-b/a =4 product of roots are c/a=1 hence roots are 1/4 and 4/2

OpenStudy (anonymous):

d(approximately equals)3.7321,0.2679

OpenStudy (radar):

\[d ^{2}-4d+1=0\] a=1 b=-4 c=1 Plug these into the quadratic formula. See electrokid's post for the formula.

OpenStudy (radar):